# On the complexity of 2-bridge link complements

### v1

@@ 1 Introduction
Complexity bounds for Dehn-fillings have been determined by Chu [5, 6, 7], and in more special cases in [17]. In [20], Lackenby and Purcell give complexity bounds for fibred manifolds based on properties of the monodromy gluing the fibres.
@@ Acknowledgements.
The authors thank Stephan Tillmann for the initial ideas that lead to this paper and for useful discussions during its development. We also thank Technische Universität Berlin and Freie Universität Berlin for hosting the authors during October and November 2023. James Morgan is supported by an Australian Government Research Training Program (RTP) Scholarship. The work of Jonathan Spreer is partially supported by the Australian Research Council under the Discovery Project scheme (grant number DP190102259).
@@ 2.3 Hyperbolic volume and angle structures
The set of all angle structures on a triangulation is denoted $\mathcal{A}(\mathcal{T})$.
@@ 3 Sakuma-Weeks triangulations of 2-bridge link complements
We now outline the construction of the Sakuma-Weeks triangulations for the complements of $2$-bridge links and detail the combinatorics needed for our results. We follow the setup described in Section 2.1. Further details of this construction can be found in [12, 27, 30]. Throughout this section we define $\Omega=R^{a_{1}}L^{a_{2}}\cdots(R^{a_{n}}\mid L^{a_{n}})$ with $K(\Omega)$ denoting the associate $2$-bridge link, where $(R^{a_{n}}\mid L^{a_{n}})$ denotes the regular expression determining either the string $R^{a_{n}}$ or $L^{a_{n}}$ which we assume is done appropriately. We assume that $a_{i}>0$ for all $1\leq i\leq n$ and that $a_{1},a_{n}\geq 1$. We set $\ell=\sum_{i=1}^{n}a_{i}$. Let $M=S^{3}\backslash K(\Omega)$ denote the link complement. It is vital to first note that we consider a 2-bridge link as being embedded in $S^{3}$; we suppress this for most of what follows.
@@ 3.1 Construction of the Sakuma-Weeks triangulations
By gluing $S_{i}\times\{0\}$ and $S_{i+1}\times\{1\}$ along their horizontal and vertical edges we obtain two regions, each bounded by four ideal vertices, six ideal edges, and four triangular faces. This forms a pair of ideal tetrahedra for each $11\leq i\leq\ell-$ 1. We call each such pair of tetrahedra a layer and denote this by $\widetilde{\Delta}_{i}$. These layers are glued together by identifying the diagonal edges in $S_{i}\times\{0\}$ to either the horizontal or vertical edges in $S_{i+1}\times\{1\}$. This process is shown in Figure 4. In order to complete the triangulation of the complement of the $2$-bridge link $K(\Omega)$ we identify the four remaining ideal triangles on $S_{1}\times\{1\}$ in pairs following the isotopy defined by the added exterior crossing, and similarly for the four remaining faces on $S_{\ell-1}\times\{0\}$. The identification on $S_{\ell-1}\times\{0\}$ is shown in Figure 5. Since we assume that the $\Omega$ starts with $R^{a_{1}}$ we can note that the ideal triangles on $S_{1}\times\{1\}$ are identified similarly to those on $S_{\ell-1}\times\{0\}$ in the case where $\Omega$ ends with $R^{a_{n}}$. The resulting triangulation $\mathcal{T}=\mathcal{T}(M)$ consists of layers $\widetilde{\Delta}_{1},\dots,\widetilde{\Delta}_{\ell-1}$, each consisting of two ideal tetrahedra $\Delta_{2i-1}$ and $\Delta_{2i}$ for $11\leq i\leq\ell-$ 1. This gives the size of the Sakuma-Weeks triangulation as $|\mathcal{T}|=2(\ell-1)=2\sum_{i=1}^{n}a_{i}-2.$
Figure 4: The process of gluing the layers of the Sakuma-Weeks triangulation together. Each layer consists of two ideal tetrahedra glued along their horizontal and vertical edges. The four faces on the back of $\widetilde{\Delta_{i}}$ lie on $S_{i}\times\{0\}$ and the four faces on the front lie on $S_{i+1}\times\{1\}$ and similarly for $\widetilde{\Delta_{i+1}}$. We have highlighted two faces on $S_{i}\times\{1\}$ and the faces they glue to in $S_{i+1}\times\{0\}$.
Figure 5: Edge identifications on $S_{\ell-1}\times\{0\}$ when $\Omega$ ends with $R$ (left) and $L$ (right). For each case, the shaded faces are identified and the unshaded faces are identified. Note that no two faces of a single ideal tetrahedron are identified.
@@ 3.2 Edge Degrees
The isotopy defined by the crossing between $S_{i}\times\{1\}$ and $S_{i}\times\{0\}$ combined with the gluing of $S_{i}\times\{0\}$ and $S_{i+1}\times\{1\}$ along vertical and horizontal edges allows us to determine each edge class of $\mathcal{T}$ explicitly and hence compute their degrees. Figure 4 shows the the tracking of edges through the triangulation with Figure 5 showing the edge identifications coming from the gluing of faces on $S_{1}\times\{1\}$ and $S_{\ell}\times\{0\}$. We make two immediate observations from these figures.
@@
If $a_{1}=1$, then the first two letters of $\Omega$ are $RL$ and we see that a vertical edges in $\widetilde{\Delta}_{1}$ are glued to vertical edges in $\widetilde{\Delta}_{2}$. If $a_{1}>1$, however, then the vertical edges in $\widetilde{\Delta}_{1}$ are glued to distinct diagonal edges in $\widetilde{\Delta}_{2}$ giving edge classes of degree 3. The argument for $a_{n}>1$ is analagous, noting that the last letter of $\Omega$ only determines how the remaining four faces are identified on the last layer. ∎
@@ 5 Complexity bounds via angle structures
We utilise this lemma to simplify the construction of explicit angle structures. In an abuse of notation, we refer to both the $i$-th layer and the tetrahedra it contains as $\Delta_{i}$. Throughout this section we assume the conditions from Theorem 11.
@@ 5.1 Block decomposition of 2-bridge links
Let $(a_{i},\dots,a_{i+k})$ be a maximal sequence of exponents such that $a_{j}=1$ for $i\leq j\leq i+k$ with $k\geq 0$. This sequence has length $k+1$. If $n=k+1$, then the sequence corresponds to a (B1) block. If $n>k+1$, $i\not\in\{1,n\}$, then $a_{i-1}=a_{i+1}=2$ by the maximality of the sequence. It follows that the sequence corresponds to the start, the middle, or the end of a (B3) block. Similarly, if $n>k+1$, $i=1,n$, then the sequence corresponds to the start or the end of a (B3) block, respectively.
@@ 5.2 Angle structures on blocks
We show that a decomposition of the Sakuma-Weeks triangulation $\mathcal{T}$ associated to a word $\Omega$ can be endowed with an angle structure $\Theta\in\mathcal{A}(\mathcal{T})$ which can be computed directly from the decomposition of $\Omega^{\prime}$ into blocks of types (B1), (B2) and (B3). Given a block of $\Omega^{\prime}$ corresponding to one of the types in Definition 13 we can find the corresponding subcomplex $\mathcal{S}\subseteq\mathcal{T}$. The edge classes which glue to adjacent subcomplexes are incomplete. From condition (2) of Definition 3, the angle sum around any edge class in $\mathcal{T}$ adds to $2\pi$. We define the angle deficit as the remaining angle required to achieve $2\pi$ on such incomplete edge classes. We call an edge an interior edge if its edge class is entirely contained in a single block, otherwise we call it a boundary edge. Each block has three boundary edge classes at both the start and end of the corresponding subcomplex, each with an angle deficit. We call the triple of angle deficits at the start or end of a block a $\delta$- or $\epsilon$-boundary deficit, respectively. Recall that $\Theta=(\theta_{1},\dots,\theta_{3\ell})\in\mathcal{A}(\mathcal{T})$ where $\ell=|\Omega|-1$ and the angles assigned to the $i$-th tetrahedron are $\Theta^{(i)}=(\theta_{3i-2},\theta_{3i-1},\theta_{3i})$. We define the boundary deficits on a block $\delta=(\delta_{0},\delta_{1},\delta_{2})$ and $\epsilon=(\epsilon_{0},\epsilon_{1},\epsilon_{2})$ in the same way (Figure 8).
Figure 8: Assignment of $\delta=(\delta_{0},\delta_{1},\delta_{2})$ and $\epsilon=(\epsilon_{0},\epsilon_{1},\epsilon_{2})$ boundary deficits on a block containing $k+1$ layers.
@@
Suppose that for $\Omega^{\prime}$ we achieve the desired boundary deficits with each tetrahedron being regular ideal for some $m\geq 1$ and $p\in\{0,1\}$. The angle equation of each interior edge of the block is either
@@
for $k\geq 0$. In each case, setting each angle to be $\pi/3$ satisfies these equations. The angle equations for boundary edge classes at the start of the block are,
@@
One (B2) block – Let $\Omega^{\prime}=L(R^{2}L^{2})^{m}(R^{2}L)^{p}R^{p-1}$ with $m\geq 0$, $p\in\{0,1\}$. When $m=0$, $p=1$ we have $\Omega^{\prime}=LR^{2}L$ giving the following angle equations,
@@
$\Theta^{(i+2)}=\left(\frac{\pi}{4},\;\frac{\pi}{4},\;\frac{\pi}{2}\right)\qquad\text{and}\qquad\Theta^{(i+3)}=\left(\frac{\pi}{2},\;\frac{\pi}{4},\;\frac{\pi}{4}\right).$ (5.6)
Observe that in the preceding two cases the first two angle vectors are the same and the last two angle vectors are the same with the first two angles swapped. The order in which these angles are assigned can be determined by the last letter of the block. If $\Omega^{\prime}$ ends on $L$, then we assign these angles as in Equation 5.4, otherwise we set them as in Equation 5.5. For this assignment of angles we always obtain the desired boundary deficits. Similarly, when increasing $m$, if the added pair of syllables is $L^{2}R^{2}$ we repeat the assignment of angles in Equation 5.6 with the first two angles swapped in each vector.
@@
for $k\geq 1$. In the case that Equation 5.7 is followed by Equation 5.8 we can solve the equations by setting the relevant tetrahedra to have shape (III) as
@@
Noting that $\Omega^{\prime}$ ends on $L$, we set $\Theta^{(i)}$, $\Theta^{(i+1)}$, $\Theta^{(i+5)}$ and $\Theta^{(i+6)}$ as in Equation 5.4 in order to obtain tetrahedra of types (I), (VI), (IV) and (II). To obtain tetrahedra of type (III) we set,
@@
Suppose that for some $m>2$ we can assign angles as above with tetrahedra shapes (I) and (VI) appearing at the start of the block, shapes (IV) and (II) appearing at the end of the block and all other tetrahedra having shapes (III) and (VIII) following the pattern of $\Theta^{(i+2)},\Theta^{(i+3)}$ and $\Theta^{(i+4)}$ above. Consider $\Omega^{\prime}=L(R^{2}L)^{q+1}$. The modified and new angle equations are
@@
By our assumption $\Delta_{(i+3m-1)}$ has shape (VIII). We set $\Delta_{i+3m+3}$ and $\Delta_{i+3m+4}$ to have shapes (IV) and (VI), respectively. Substituting these values into the above equations we find
@@
where $\mu_{0},\mu_{1},\mu_{2}$ denote the angle deficits on $\mathcal{S}^{\prime\prime}$, replacing the angles from $\Theta^{(k)}$ for $k\geq 1+3$. Since $\delta_{0}=\pi/3$ we solve the above equations as
@@
Suppose we now extend the length of the (B2) block in $\Omega^{\prime\prime}$ by one. This corresponds to a subword of $\Omega^{\prime}=(LR^{2}L^{2})(RL^{2})^{q}R$ with $q=m-1$. By assumption we have $\Theta^{(i)},\Theta^{(i+1)}$ and $\Theta^{(i+2)}$ as above and further that $\Theta^{(i+3)}$ and $\Theta^{(i+4)}$ have shape (III). The relevant equations containing the new deficits of $\mathcal{S}^{\prime\prime}$ are lines 2, 4, and 5 of Equation 5.12. Substituting the angles in we have
@@
where $\eta_{0},\eta_{1},\eta_{2}$ denote the new deficits on $\mathcal{S}^{\prime\prime}$, replacing the angles from $\Theta^{(k)}$ for $k\geq 3m+4$. Solving the above equations we see that
@@
Since $\mu_{i}=\eta_{i}$ for each $i=1,2,3$, extending any (B2) block in $\Omega^{\prime}$ must introduce ideal tetrahedra with shape (III). The remaining shapes of the tetrahedra in both instances remain the same. However, after we extend the (B2) block we must swap the first two angles in each remaining $\Theta^{(j)}$ as described in Cases 1 and 2 above. We also observe that the angle deficits seen here are the same if we follow this procedure at the end of any (B2) block in $\Omega^{\prime}$. ∎
@@
Denote the first and last letters of a (B1) and (B3) block by $\alpha_{(B1)}$, $\alpha_{(B3)}$, and $\omega_{(B1)}$, $\omega_{(B3)}$, respectively. Let $\epsilon_{0}$, $\epsilon_{1}$ and $\epsilon_{2}$ denote the boundary deficits at the end of the (B1) block as in Figure 8. Similarly, let $\delta_{0}$, $\delta_{1}$ and $\delta_{2}$ be the boundary deficits at the start of the (B3) block. One key take away from Lemmas 15 and 17 is that the $\epsilon$-boundary deficit of a (B1) block equals the $\delta$-boundary deficit of a block of (B3) as long as $\omega_{(B1)}\neq\alpha_{(B3)}$. In particular, we have $\delta_{i}=\epsilon_{i}$, $1\leq i\leq 3$ whenever $\omega_{(B1)}\neq\alpha_{(B3)}$. The same is true for the end of a (B3) block and the start of a (B1) block.
@@
From Figure 5 we equate this boundary deficit to edges in $\Delta_{1}$ as,
@@
We now consider the shape of $\Delta_{\Sigma}$. Suppoe that $\Omega=R\cdots R^{a_{n-1}}L^{a_{n}}R$. From Figures 4 and 5 the angle equations for edges in $\Delta_{\Sigma}$ are
@@
where $x$ denotes the angle deficit on the vertical edge class of $\Delta_{\Sigma}$. We consider two cases for the values of $a_{n-1}$ and $a_{n}$.
@@
where $x_{0}$, $x_{1}$ and $x_{2}$ denote angle deficits on the respective edge classes. For any angle structure on this block to be compatible with a (B1) or (B3) block we require that the $\epsilon$-boundary deficit on $\Delta_{4m-1}$ agrees with the $\delta$-boundary deficit of the (B1) or (B3) blocks - which must start on $R$. That is $\epsilon_{i}=\delta_{i}$ for $1\leq i\leq 3$. Hence $\epsilon_{0}=\pi$, $\epsilon_{1}=\pi/3$ and $\epsilon_{2}=5\pi/3$.
@@
Suppose that $m=1$ so that the (B2) block is of length one. The angle deficits above must be satisfied by the angle equations from $\Delta_{1}$. From Figures 4 and 5 we have,
@@
We have $\pi/2=x_{0}=\theta_{3(4m-3)}$ and assume $\Delta_{4m-3}$ has shape (III), hence we assign $\Theta^{(4m-3)}=(\pi/4,\pi/4,\pi/2)$. By construction, $y$ must correspond to the horizontal edge class of the previous layer. Observe that $\Omega^{\prime}$ repeats strings of $(L^{2}R^{2})$. This corresponds to Equation 5.8 followed by Equation 5.7 and hence we have $\Theta^{(4m-4)}=(\pi/4,\pi/2,\pi/4)$. This forces $y=3\pi/4$. From the proof of lemma 17 we know that $y=2\pi/4+\pi/4$ and thus conclude that $\Delta_{4m-4}$ and $\Delta_{4m-3}$ have shape (III).

### v2

@@ 1 Introduction
Complexity bounds for Dehn-fillings have been determined by Cha [5, 6, 7], and in more special cases in [17]. In [20], Lackenby and Purcell give complexity bounds for fibred manifolds based on properties of the monodromy gluing the fibres.
@@ Acknowledgements.
The authors thank Stephan Tillmann for the initial ideas that lead to this paper and for useful discussions during its development. The authors also thank the anonymous referee for many helpful comments that improved this paper. We thank Technische Universität Berlin and Freie Universität Berlin for hosting the authors during October and November 2023. The work of both authors is partially supported by the Australian Research Council under the Discovery Project scheme (grant number DP190102259). James Morgan is additionally supported by an Australian Government Research Training Program (RTP) Scholarship.
@@ 2.3 Hyperbolic volume and angle structures
The set of all angle structures on a triangulation is denoted $\mathcal{A}(\mathcal{T})$. The assignment of dihedral angles to the $i$–th ideal tetrahedron is notated as $\Theta^{(i)}=(\theta_{3i-2},\theta_{3i-1},\theta_{3i})$.
@@ 3 Sakuma-Weeks triangulations of 2-bridge link complements
We now outline the construction of the Sakuma-Weeks triangulations for the complements of $2$-bridge links and detail the combinatorics needed for our results. We follow the setup described in Section 2.1. Further details of this construction can be found in [12, 27, 30]. Throughout this section we define $\Omega=R^{a_{1}}L^{a_{2}}\cdots(R^{a_{n}}\mid L^{a_{n}})$ with $K(\Omega)$ denoting the associated $2$-bridge link, where $(R^{a_{n}}\mid L^{a_{n}})$ denotes the regular expression determining either the string $R^{a_{n}}$ or $L^{a_{n}}$ which we assume is done appropriately. We assume that $a_{i}>0$ for all $1\leq i\leq n$ and that $a_{1},a_{n}\geq 1$. We set $\ell=\sum_{i=1}^{n}a_{i}$. Let $M=S^{3}\backslash K(\Omega)$ denote the link complement. It is vital to first note that we consider a 2-bridge link as being embedded in $S^{3}$; we suppress this for most of what follows.
@@ 3.1 Construction of the Sakuma-Weeks triangulations
By gluing $S_{i}\times\{0\}$ and $S_{i+1}\times\{1\}$ along their horizontal and vertical edges we obtain two regions, each bounded by four ideal vertices, six ideal edges, and four triangular faces. This forms a pair of ideal tetrahedra for each $11\leq i\leq\ell-$ 1. We call each such pair of tetrahedra a layer and denote this by $\widetilde{\Delta}_{i}$. These layers are glued together by identifying the diagonal edges in $S_{i+1}\times\{1\}$ to either the horizontal or vertical edges in $S_{i+1}\times\{0\}$ for $21\leq i\leq\ell-$ 2. This process is shown in Figure 4. There remains four unidentified faces in $\widetilde{\Delta}_{1}$ and four unidentified faces in $\widetilde{\Delta}_{\ell-1}$ coming from $S_{1}\times\{0\}$ and $S_{\ell}\times\{1\}$, respectively. In order to complete the triangulation of the complement of the $2$-bridge link $K(\Omega)$ we identify the four remaining ideal triangles on $S_{1}\times\{0\}$ in pairs following the isotopy defined by the crossing in $S_{1}$ and the added exterior crossing. Similarly, for the four remaining faces on $S_{\ell}\times\{1\}$, we follow the isotopy defined by the crossing in $S_{\ell}$ and the added interior crossing. The identification on $S_{\ell-1}\times\{0\}$ is shown in Figure 5. Since we assume that the $\Omega$ starts with $R^{a_{1}}$ we can note that the ideal triangles on $S_{1}\times\{1\}$ are identified similarly to those on $S_{\ell-1}\times\{0\}$ in the case where $\Omega$ ends with $R^{a_{n}}$. The resulting triangulation $\mathcal{T}=\mathcal{T}(M)$ consists of layers $\widetilde{\Delta}_{1},\dots,\widetilde{\Delta}_{\ell-1}$, each consisting of two ideal tetrahedra $\Delta_{2i-1}$ and $\Delta_{2i}$ for $11\leq i\leq\ell-$ 1. This gives the size of the Sakuma-Weeks triangulation as $|\mathcal{T}|=2(\ell-1)=2\sum_{i=1}^{n}a_{i}-2.$
Figure 4: The process of gluing the layers of the Sakuma-Weeks triangulation together. Each layer consists of two ideal tetrahedra glued along their horizontal and vertical edges. The four faces on the back of $\widetilde{\Delta_{i}}$ lie on $S_{i}\times\{0\}$ and the four faces on the front lie on $S_{i+1}\times\{1\}$ and similarly for $\widetilde{\Delta_{i+1}}$. We have highlighted two faces on $S_{i+1}\times\{1\}$ and the faces they glue to in $S_{i+1}\times\{0\}$.
(a) Identifications of ideal triangles on $S_{1}\times\{0\}$. The shaded faces are identified and the unshaded faces are identified.
(b) Identifications of ideal triangles on $S_{\ell}\times\{1\}$ when $\Omega$ ends with $R$ (left) and $L$ (right). For each case, the shaded faces are identified and the unshaded faces are identified.
Figure 5: Identifications of remaining ideal triangles in $\widetilde{\Delta}_{1}$ and $\widetilde{\Delta}_{2}$. For each layer shown, the back faces belong to $S_{i}\times\{0\}$ and the front faces belong to $S_{i}\times\{1\}$. Note that no two faces of a single ideal tetrahedron are identified.
@@ 3.2 Edge Degrees
The isotopy defined by the crossing between $S_{i}\times\{1\}$ and $S_{i}\times\{0\}$ combined with the gluing of $S_{i}\times\{0\}$ and $S_{i+1}\times\{1\}$ along vertical and horizontal edges allows us to determine each edge class of $\mathcal{T}$ explicitly and hence compute their degrees. Figure 4 shows the tracking of edges through the triangulation with Figure 5 showing the edge identifications coming from the gluing of faces on $S_{1}\times\{0\}$ (Figure 5(a)) and $S_{\ell}\times\{1\}$ (Figure 5(b)). We make two immediate observations from these figures.
@@
If $a_{1}=1$, then the first two letters of $\Omega$ are $RL$ and we see that the vertical edges in $\widetilde{\Delta}_{1}$ are glued to vertical edges in $\widetilde{\Delta}_{2}$. If $a_{1}>1$, however, then the vertical edges in $\widetilde{\Delta}_{1}$ are glued to distinct diagonal edges in $\widetilde{\Delta}_{2}$ giving edge classes of degree 3. The argument for $a_{n}>1$ is analagous, noting that the last letter of $\Omega$ only determines how the remaining four faces are identified on the last layer. ∎
@@ 5 Complexity bounds via angle structures
We utilise this lemma to simplify the construction of explicit angle structures. For the remainder of this paper, in an abuse of notation, we refer to both the $i$-th layer and the tetrahedra it contains as $\Delta_{i}$. Throughout this section we assume the conditions from Theorem 11.
@@ 5.1 Block decomposition of 2-bridge links
Let $(a_{i},\dots,a_{i+k})$ be a maximal sequence of exponents such that $a_{j}=1$ for $i\leq j\leq i+k$ with $k\geq 0$. This sequence has length $k+1$. If $n=k+1$, then $i=1$ and the sequence corresponds to a (B1) block. Hence let $n>k+1$. If $i=1$, then $a_{i+k+1}=2$ and the (possibly empty) subsequence $(a_{1},\dots,a_{i+k-1})$ corresponds to a (B1) block with $(a_{i+k},a_{i+k+1})$ corresponding to the start of a (B3) block. Similarly, if $i+k=n$, then $a_{i-1}=2$ and $(a_{i-1},a_{i})$ corresponds to the end of a (B3) block followed by a (possibly empty) (B1) block. Otherwise $a_{i-1}=a_{i+k+1}=2$. If $k=0$, then the sequence is contained in a (B3) block; if $k\geqq 1$, then the sequence corresponds to the end of a (B3) block followed by a (possibly empty) (B1) block followed by the start of another (B3) block.
@@ 5.2 Angle structures on blocks
We show that a decomposition of the Sakuma-Weeks triangulation $\mathcal{T}$ associated to a word $\Omega$ can be endowed with an angle structure $\Theta\in\mathcal{A}(\mathcal{T})$ which can be computed directly from the decomposition of $\Omega^{\prime}$ into blocks of types (B1), (B2) and (B3). Given a block of $\Omega^{\prime}$ corresponding to one of the types in Definition 13 we can find the corresponding subcomplex $\mathcal{S}\subseteq\mathcal{T}$. The edge classes which glue to adjacent subcomplexes are incomplete. From condition (2) of Definition 3, the angle sum around any edge class in $\mathcal{T}$ adds to $2\pi$. We define the angle deficit as the remaining angle required to achieve $2\pi$ on such incomplete edge classes. We call an edge an interior edge if its edge class is entirely contained in a single block, otherwise we call it a boundary edge. Each block has three boundary edge classes at both the start and end of the corresponding subcomplex, each with an angle deficit. We call the triple of angle deficits at the start or end of a block a $\delta$- or $\epsilon$-boundary deficit, respectively. Recall from Definition 3 that $\Theta=(\theta_{1},\dots,\theta_{3\ell})\in\mathcal{A}(\mathcal{T})$ where $\ell=|\Omega|-1$ and the angles assigned to the $i$-th tetrahedron are $\Theta^{(i)}=(\theta_{3i-2},\theta_{3i-1},\theta_{3i})$. We define the boundary deficits on a block $\delta=(\delta_{0},\delta_{1},\delta_{2})$ and $\epsilon=(\epsilon_{0},\epsilon_{1},\epsilon_{2})$ in the same way (Figure 8).
Figure 8: Assignment of $\delta=(\delta_{0},\delta_{1},\delta_{2})$ and $\epsilon=(\epsilon_{0},\epsilon_{1},\epsilon_{2})$ boundary deficits on a block containing $k+1$ layers, starting from layer $\Delta_{i}$ and containing all layers up to (and including) $\Delta_{i+k}$.
@@
Suppose that for $\Omega^{\prime}$ we achieve the desired boundary deficits with each tetrahedron being regular ideal for some $m\geq 1$ and $p\in\{0,1\}$. The angle equation an interior edge of the block first appearing in $\Delta_{i+k}$, $k\geq 0$, is either
@@
In each case, setting each angle to be $\pi/3$ satisfies these equations. The angle equations for boundary edge classes at the start of the block are,
@@
One (B2) block – Let $\Omega^{\prime}=L(R^{2}L^{2})^{m}(R^{2}L)^{p}R^{1-p}$ with $m\geq 0$, $p\in\{0,1\}$. When $m=0$, $p=1$ we have $\Omega^{\prime}=LR^{2}L$ giving the following angle equations,
@@
$\Theta^{(i+2)}=\left(\frac{\pi}{4},\;\frac{\pi}{4},\;\frac{\pi}{2}\right)\qquad\text{and}\qquad\Theta^{(i+3)}=\left(\frac{\pi}{2},\;\frac{\pi}{4},\;\frac{\pi}{4}\right),$ (5.6)
yielding $\epsilon_{0}=\pi/3$. Observe that in the preceding two cases the first two angle vectors are the same and the last two angle vectors are the same with the first two angles swapped. The order in which these angles are assigned can be determined by the last letter of the block. If $\Omega^{\prime}$ ends on $L$, then we assign these angles as in Equation 5.4, otherwise we set them as in Equation 5.5. For this assignment of angles we always obtain the desired boundary deficits. Similarly, when increasing $m$, if the added pair of syllables is $L^{2}R^{2}$ we repeat the assignment of angles in Equation 5.6 with the first two angles swapped in each vector.
@@
for $k\geq 1$. In the case that Equation 5.7 is followed by Equation 5.8, obtained by replacing $(i+k)$ in the latter equation with $(i+k+2)$, we can solve the equations by setting the relevant tetrahedra to have shape (III) as
@@
Noting that $\Omega^{\prime}$ ends on $L$, we set $\Theta^{(i)}$, $\Theta^{(i+1)}$, $\Theta^{(i+5)}$ and $\Theta^{(i+6)}$ as in Equation 5.4 in order to obtain tetrahedra of types (I), (VI), (IV) and (II), respectively. To obtain tetrahedra of type (III) we set,
@@
Suppose that for some $m>2$ we can assign angles as above with tetrahedra shapes (I) and (VI) appearing at the start of the block, shapes (IV) and (II) appearing at the end of the block and all other tetrahedra having shapes (III) and (VIII) following the pattern of $\Theta^{(i+2)},\Theta^{(i+3)}$ and $\Theta^{(i+4)}$ above. Consider $\Omega^{\prime}=L(R^{2}L)^{m+1}$. The modified and new angle equations are
@@
By our assumption $\Delta_{i+3m-1}$ has shape (VIII). We set $\Delta_{i+3m+3}$ and $\Delta_{i+3m+4}$ to have shapes (IV) and (II), respectively, as in Equation 5.4. Substituting these values into the above equations we find
@@
where $\mu_{0},\mu_{1},\mu_{2}$ denote the angle deficits on $\mathcal{S}^{\prime\prime}$, replacing the angles from $\Theta^{(k)}$ for $k\geq i+3$ in Equation 5.9. Since $\delta_{0}=\pi/3$ we solve the above equations as
@@
Suppose we now extend the length of the (B2) block in $\Omega^{\prime\prime}$ by one. This corresponds to a subword of $\Omega^{\prime}=(LR^{2}L^{2})(RL^{2})^{q}R$ with $q=m-1$. By assumption we have $\Theta^{(i)}$ and $\Theta^{(i+1)}$ as above and further that $\Theta^{(i+2)},\Theta^{(i+3)}$ and $\Theta^{(i+4)}$ have shape (III) assigned as
$\Theta^{(i+2)}=\left(\frac{\pi}{4},\frac{\pi}{4},\frac{\pi}{2}\right),\quad\Theta^{(i+3)}=\left(\frac{\pi}{2},\frac{\pi}{4},\frac{\pi}{4}\right),\quad\text{and}\quad\Theta^{(i+4)}=\left(\frac{\pi}{4},\frac{\pi}{4},\frac{\pi}{2}\right).$
The relevant equations containing the new deficits of $\mathcal{S}^{\prime\prime}$ are
$\theta_{3(i+1)}+2\theta_{3(i+2)-1}+2\theta_{3(i+3)-1}+2\theta_{3(i+4)-1}+\eta_{0}$ $=2\pi,$
$\theta_{3(i+3)}+2\theta_{3(i+4)-2}+\eta_{1}$ $=2\pi,$
$\theta_{3(i+4)}+\eta_{2}$ $=2\pi,$
where $\eta_{0},\eta_{1},\eta_{2}$ denote the new deficits on $\mathcal{S}^{\prime\prime}$. Substituting the angles in we have
@@
Solving the above equations we see that
@@
Since $\mu_{i}=\eta_{i}$ for each $i=0,1,2$, extending any (B2) block in $\Omega^{\prime}$ must introduce ideal tetrahedra with shape (III). The remaining shapes of the tetrahedra in both instances remain the same. However, after we extend the (B2) block we must swap the first two angles in each remaining $\Theta^{(j)}$ as described in Cases 1 and 2 above. We also observe that the angle deficits seen here are the same if we follow this procedure at the end of any (B2) block in $\Omega^{\prime}$. ∎
@@
Denote the first and last letters of a (B1) and (B3) block by $\alpha_{(B1)}$, $\alpha_{(B3)}$, and $\omega_{(B1)}$, $\omega_{(B3)}$, respectively. Let $\epsilon_{0}$, $\epsilon_{1}$ and $\epsilon_{2}$ denote the boundary deficits at the end of the (B1) block as in Figure 8. Similarly, let $\delta_{0}$, $\delta_{1}$ and $\delta_{2}$ be the boundary deficits at the start of the (B3) block. One key take away from Lemmas 15 and 17 is that the $\epsilon$-boundary deficit of a (B1) block equals the $\delta$-boundary deficit of a block of (B3) as long as $\omega_{(B1)}\neq\alpha_{(B3)}$. In particular, we have $\delta_{i}=\epsilon_{i}$, $0\leq i\leq 2$ whenever $\omega_{(B1)}\neq\alpha_{(B3)}$. The same is true for the end of a (B3) block and the start of a (B1) block.
@@
From Figure 5(a) we equate this boundary deficit to edges in $\Delta_{1}$ as,
@@
We now consider the shape of $\Delta_{\Sigma}$. Suppose that $\Omega=R\cdots R^{a_{n-1}}L^{a_{n}}R$ so that $n$ is odd. From Figures 4 and 5(b) (using the case of $n$ odd), the angle equations for edges in $\Delta_{\Sigma}$ are
@@
where $x$ denotes the angle deficit on the vertical edge class of $\Delta_{\Sigma}$. To verify this, we note that the four horizontal edges and two diagonal edges in $\Delta_{\Sigma}$ are identified and, given $\Omega$, must also include both diagonal edges from the previous layer. This gives the first equation. To obtain the second equation, note that we have two distinct edge classes for the four vertical edges in $\Delta_{\Sigma}$ which also include some edges from previous layers. We consider two cases for the values of $a_{n-1}$ and $a_{n}$.
@@
where $x_{0}$, $x_{1}$ and $x_{2}$ denote angle deficits on the respective edge classes. For any angle structure on this block to be compatible with a (B1) or (B3) block we require that the $\epsilon$-boundary deficit on $\Delta_{4m-1}$ agrees with the $\delta$-boundary deficit of the (B1) or (B3) blocks - which must start on $R$. That is $\epsilon_{i}=\delta_{i}$ for $0\leq i\leq 2$. Hence $\epsilon_{0}=\pi$, $\epsilon_{1}=\pi/3$ and $\epsilon_{2}=5\pi/3$.
@@
Suppose that $m=1$ so that the (B2) block is of length one. The angle deficits above must be satisfied by the angle equations from $\Delta_{1}$. From Figures 4 and 5(a) we have,
@@
We have $\pi/2=x_{0}=\theta_{3(4m-3)}$ and assume $\Delta_{4m-3}$ has shape (III), hence we assign $\Theta^{(4m-3)}=(\pi/4,\pi/4,\pi/2)$. By construction, $y$ must correspond to the horizontal edge class of the previous layer. Observe that $\Omega^{\prime}$ repeats strings of $(L^{2}R^{2})$. This corresponds to Equation 5.8 followed by Equation 5.7 and hence, from the reasoning following these equations, we have $\Theta^{(4m-4)}=(\pi/4,\pi/2,\pi/4)$. This forces $y=3\pi/4$. From the proof of Lemma 17 (the case of one (B2) block) we know that $y=2\pi/4+\pi/4$ and thus conclude that $\Delta_{4m-4}$ and $\Delta_{4m-3}$ have shape (III).